OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 2x - cos 2x 2 sin2x - 2 sin x cos x + 1 2 sin x 2 sin2x + 2 sin x cos x - 1 2 sin2x - 2 sin x cos x - 1

OpenStudy (whpalmer4):

double angle formulas: \[\sin 2u = 2 \sin u \cos u\]\[\cos 2u = \cos^2 u - \sin^2u = 2 \cos^2u-1 = 1-2\sin^2u\]